# Advancing Galaxy Analysis: AI-Powered Detection and Segmentation of Edge-On Galaxies

### v1

@@ 5 Conclusions
We present a novel deep learning algorithm for the detection and extraction of edge-on galaxies in astronomical images. Based on Galaxy Zoo classification, we obtained images of approximately $16,000$ edge-on, spiral galaxies from the SDSS catalogue, that we trained the algorithm on. Our results show that the algorithm can reliably detect edge-on galaxies, with Precision=$0.80$ and Recall=$0.94$. We also successfully applied a segmentation algorithm to extract the galaxies from the image, so they can be used for future studies. We retrieved the parameters of our sample of galaxies from the SDSS and provide them together with confidence values. From a basic analysis of the properties of our sample, we find that the majority of galaxies is nearby ($0.100.02<z<$ 0.10), red and have low values of $b/a$, which is in agreement with other works. Our analysis of the SDSS data showed that we can reliably detect the needed galaxies, however, the true potential of the algorithm lies in its possible application on data from future surveys, such as Euclid (Euclid Collaboration et al., 2024), Roman Space Telescope (Bartusek et al., 2022) or Large Synoptic Survey Telescope (LSST, Ivezić et al., 2019) where we expect a significant increase in data volume, which necessitates processing through machine learning.
Acknowledgements
@@
Data Availability
The catalogue of galaxies generated in this work is available at our Github repository https://github.com/MatFyzGalaxy/EdgeOnML.

### v2

@@ 5 Conclusions
We present a novel deep learning algorithm for the detection and extraction of edge-on galaxies in astronomical images. Based on Galaxy Zoo classification, we obtained images of nearly $16,000$ edge-on, spiral galaxies from the SDSS catalog that we trained the algorithm on. Our results show that the algorithm can reliably detect edge-on galaxies, with Precision=0.80 and Recall=0.94. We also successfully applied a segmentation algorithm to extract the galaxies from the image, so they can be used for future studies. We retrieved the parameters of our sample of galaxies from the SDSS and provide them together with confidence values. From a basic analysis of the properties of our sample, we find that the majority of galaxies are nearby ($0.100.02<z<$ 0.10), red and have low values of $b/a$, which is in agreement with other works, as well as the training data. Our analysis of the SDSS data showed that we can reliably detect the needed galaxies, however, the true potential of the algorithm lies in its possible application on data from future surveys, such as Euclid (Euclid Collaboration et al., 2024), Roman Space Telescope (Bartusek et al., 2022) or Large Synoptic Survey Telescope (LSST, Ivezić et al., 2019) where we expect a significant increase in data volume, which necessitates processing through machine learning. In a future work, we will develop this algorithm to automatically detect warped galaxies.